the area of a plot of land is 126 square meter.If the breadth of plot is 5 meter less then its lenght,what is its length and breadth
Answer by ikleyn(52781)

.
W*(W+5) = 126.


Solve for W and then evaluate L = W + 5.


Or solve it mentally  126 = 9*14  is decomposition of the number 126 into the product of two positive numbers with the difference of 5.